Two alternative designs are under consideration for a new ride called the Scream Machine at a theme park located in Florida. The two candidate designs differ in complexity, cost and predicted revenue. Alternative A will require an investment of $300,000 and is estimated to produce after-tax revenue of $55,000 annually over a 10-year horizon. Alternative B will require an investment of $450,000 and is expected to generate annual after-tax revenue of $80,000. A negligible salvage value is assumed for both designs. The theme park management could decide to “do nothing;” if so, the present worth of doing nothing will be zero. Which alternative design, if either, should the theme park select if its !"## = 8%?
Answer #1

NPW of first option = -300000 + 55000 * (P/A, 8%,10)

= -300000 + 55000 * 6.710081

= 69054.48

NPW of second option = -450000 + 80000 * (P/A, 8%,10)

= -450000 + 80000 * 6.710081

= 86806.51

As NPW of second option is more it should be selected